A right circular cylinder has a surface area of 112π. If the height of the cylinder is 10 find the radius of its base.
Answer by Greenfinch(383)

Surface area is top and bottom 2 x r^2 x pi
side is 2 x r x h x pi
We are told that the area is 112 x pi
so we can remove pi from the question and get
(2 x r^2) + (2 x r x h) = 112
and since h = 10 and there are factors of 2 on both sides
r^2 + 10 x r = 56
0r r^2 + 10 x r - 56 = 0
or (r - 4)(r + 14) = 0
r = 4 or -14
As -14 isn't an option, r = 4